# PsyDI: Towards a Personalized and Progressively In-depth Chatbot for Psychological Measurements

### v1

@@ Appendix A Dataset Details
The datasets used in PsyDI are described in Table 4. Each dataset is detailed with its language, the number of training and testing samples, and its origin.
@@
Reddit English 166580 1850 Reddit posts and the corresponding tags
@@
Diamonte, enerated through two ChatGPT instances
@@
Appendix B Training Details
@@
Multi-turn Chat Generation We utilize the latest version of GPT-3.5 for generating multi-turn chat interactions, as it supports JSON format output. The temperature is set to 0.7, and the top-p is set to 0.95 to ensure flexible and varied question generation.
@@ Appendix C Multi-turn Chat Prompt
Multi-turn Chat Prompt # **MBTI Cognitive Analysis Consultant** I am an MBTI cognitive analysis consultant, specializing in guiding conversations and uncovering the underlying cognitive processes of communication partners to accurately predict their MBTI types. ## Background I excel at interpreting behaviors in materials, analyzing details, and exploring the logic and cognitive patterns behind behaviors to help individuals understand their underlying cognitive processes and determine their MBTI types. ## Preferences As an MBTI cognitive analysis consultant, I value logic and clarity, preferring to use concise and approachable expressions. Additionally, I respect the viewpoints and ideas of communication partners, striving to understand their internal cognitive processes through their expressions. Furthermore, I am proficient in using Jung’s eight cognitive functions theory to comprehensively analyze individuals’ cognitive processes, understanding the specific meanings of te, ti, fe, fi, se, si, ne, and ni functions and their various potential operations in specific scenarios. I am adept at discerning MBTI personality types and skilled at judging individuals’ MBTI types based on their thinking styles and expressions, as well as summarizing information related to their MBTI types.
@@
Multi-turn Chat Prompt (Con.) ## Goals - Interpret facts in materials, analyze details, and seek possible clues related to Jung’s eight functions. - Identify meaningful and valuable details in materials regarding cognitive thinking, propose various possible behavioral manifestations resulting from different cognitive processes to help communication partners explore their underlying cognitive processes. - In the final summary, explain the user’s cognitive preferences based on Jung’s eight functions and output the MBTI type most likely to belong to the communication partner. ## Constraints - When analyzing problems, consider the elements of facts, analysis, and actions. Ask only one question at a time, inquiring about motivations or behaviors specific to the previous context. - Answer each question with options A, B, C, or D. - Present options in the first-person perspective of role-playing, representing different behavioral manifestations possibly resulting from different cognitive processes. Ensure vivid and distinct differences between options, emphasizing different cognitive directions. - There must be logical progression between multiple rounds of questioning. Your questions must differ in focus from previous rounds, concentrating on key points revealed in the communication partner’s answers, with no repetition of questions. - Always remember your role as an MBTI consultant when asking questions. Your responses must include speculation about the partner’s underlying motivations and transition to the next question. Enclose speculated motivations with *, and transitions to the next question with **, followed by a space. - Ask questions for 2-4 rounds, then directly provide the communication partner’s MBTI type and organize the conversation into the partner’s self-description, combining each question and its corresponding answer without omission. - Do not exceed four rounds of questioning. Once questioning is complete, immediately output the result following the specified format in []. - Do not directly reference questions from the Jung’s eight functions test. Questions must not contain any direct mention of MBTI. Focus solely on assisting the communication partner in exploring themselves, emphasizing key points from their answers. - If the communication partner introduces topics beyond the scope of discussion, kindly remind them to return to the conversation and proceed with questioning. ## Skills - Provide reasonable possibilities of motivation by identifying facts and analyzing answers. - Communicate viewpoints clearly and concisely. - Demonstrate strong logic, coherence in questioning, and organic expansion of topics. - Proficient in psychology, MBTI, and the interpretations of various functions in Jung’s eight cognitive functions theory.
@@ Appendix D PsyDI Online
Analysis Stage. Following multiple iterations of multi-turn questioning, PsyDI Online generates the final MBTI prediction using a multi-modal approach based on the updated MBTI profile. The final output comprises several components: the trajectory of the MBTI profile changes throughout the PsyDI test, a comprehensive analysis of the underlying cognitive functions inferred from the user’s responses, and a gemerated figure of the user’s temperament. This detailed analysis to elucidate the cognitive processes driving their behavior based on the user’s answers, while the MBTI figure visually encapsulates their personality traits.
@@ E.1 Semantic Scoring
E.2 MBTI Expression Tendencies
In this section, we provide the complete word representation to show the word tendencies of users under different MBTI. To directly show the trait of each dimension, we filter those words commonly used in both two opposite dimension and draw the other words’ frequency in Figure 12. As shown in the wordcloud, the word appears more frequent aligns with the chatacteristic of each MBTI dimension. For example, S dimension represents for sensing, which refers to a preference to pay attention to the most immediate impressions, like the "raw data" that they can see, hear and touch. Therefore, words like "activity" is the frequent words of S dimension. On the contrary, N dimension, which represents for a tendency to think about abstract things, frequently express words like "thought".
Figure 12: Word frequency of different MBTI dimension on PsyDI Online
E.3 PsyDI Pipeline Evaluation
In this section, we evaluate the effectiveness of the PsyDI pipeline from two perspectives. Firstly, the PsyDI system initializes the MBTI profile based on the user’s self-report. However, the statements in these self-reports can vary significantly. Thus, PsyDI must be able to progressively determine the user’s true MBTI profile through subsequent updates, regardless of the reliability of the initial statements.
Figure 13: MBTI rank of 16 MBTI bot under three MBTI Testing.
Figure 14: INFP Score with random initialization of MBTI profile.
To test this capability, we employed a randomized initialization of the MBTI profile and used it as a basis for further testing with an INFP bot in character AI. The aim was to observe whether PsyDI could correct the random initial results. The results of this experiment are shown in Figure 14. As depicted, the initial random statements resulted in INFP prediction scores ranging from 35 to 50. However, as the testing progressed, these scores gradually increased, eventually stabilizing above 80. This demonstrates that PsyDI accurately determine the MBTI profile through in-depth conversation in pipeline, even when starting from a poor initial state.
Subsequently, we conducted a comprehensive comparison between PsyDI and several mainstream MBTI tests in terms of their accuracy in identifying typical MBTI profiles. 16Personalities and APESK [^6] https://www.apesk.com/ are two widely recognized MBTI tests. We utilized the 16 most frequently used MBTI bots on characterAI and had each bot respond to the questions from these two tests. We recorded the rank of their actual MBTI type among all other MBTI types based on their scores. Additionally, we used PsyDI to test these 16 MBTI bots and recorded the MBTI ranking for each bot.
The comparison results are illustrated in Figure 14, where the horizontal axis represents the MBTI bot, with each corresponding point indicating the ranking of the bot’s actual MBTI within the 16 MBTI types under a specific MBTI test. For instance, INTP ranks second in the MBTI results obtained from PsyDI and APESK with INTP bot, but first in the results from 16Personalities.
As shown in the result, the average ranking accuracy of PsyDI is comparable to that of the two established tests. Moreover, for most bots, their actual MBTI type is within the top three predictions made by PsyDI. Given that the statements from typical MBTI bots are already quite representative, the effect of PsyDI’s multi-round in-depth dialogues is less pronounced. Nonetheless, these results demonstrate that PsyDI achieves an accuracy level similar to current popular MBTI tests while offering customization and personalization features.

### v3

@@ B.1 Datasets
Details of the above datasets are described in Table 4. Each dataset is detailed with its language, the number of training and testing samples, and its origin.
@@
Reddit English 166580 1850 Reddit statements and the corresponding tags
@@
Diamonte, enerated through two LLM instances
@@
B.2 Baselines
We compared the proposed score model in PsyDI with 7 existing closed-source LLMs and 3 open-source LLMs. The closed-source LLMs include GPT-3.5 Turbo and GPT-4 (Achiam et al., 2023), the acclaimed leader in LLMs, as well as DeepSeek (DeepSeek-AI, 2024), a Mixture-of-Experts (MOE) Language Model for both English and Chinese. We also evaluate the performance of Moonshot citepzhang2024moonshot, Qwen (Bai et al., 2023), Baichuan (Yang et al., 2023), and yi (Young et al., 2024). For open-source LLMs, we use Llama2 (Touvron et al., 2023), a commonly used English LLM, Chatglm3 (Du et al., 2022), an open bilingual dialogue language model, and InternLM2 (Cai & Cao, 2024).
B.3 Training Setting
We train PsyDI using llama-2-7b-hf as a base model and finetune with LoRA (Hu et al., 2021). We employ the llama-2-7b-hf model as our foundational architecture and subsequently refine it through fine-tuning with the Low-Rank Adaptation (LoRA) (Hu et al., 2021) technique. For the RL component of our methodology, we leverage the DI-engine (Niu et al., 2021) framework as our training infrastructure. Since llama-2-7b-hf itself is an English LLM, we first inject English MBTI datasets from Reddit to enhance its MBTI indicativeness judgment ability in English text, forming PsyDI-en. Then, we apply cross-lingual transfer learning (Schuster et al., 2018), finetuning PsyDI-en with both Reddit and Chinese datasets Diamente, so that PsyDI-ch retains its understanding of MBTI in English text while also mastering Chinese text.
@@ B.3 Training Setting
Multi-turn Chat Generation We utilize the latest version of GPT-3.5 for generating multi-turn chat interactions, as it supports JSON format output. The temperature is set to 0.7, and the top-p is set to 0.95 to ensure flexible and varied question generation. The length of action $L$is defined by the hyperparameter max_token, which is set to 2048.
@@
B.4 Evaluating Method
. To evaluate the score model in PsyDI, for each pair-wise data sample $(p_{i},p_{j},mbti)$, we asked all baseline LLMs to determine which statement was more likely from a user with the specified MBTI type $mbti$. The prompt was designed as follows:
All LLMs had their temperature set to 0.3 to ensure more deterministic outputs. For PsyDI, we calculate the score $-m^{\top}\cdot\sum\mathcal{F}_{m}(p)$ for each statement. Then we calculate the accuracy of LLMs in recognizing the more appropriate statement as a metric of performance.
@@ Appendix C Multi-turn Chat Prompt
Multi-turn Chat Prompt # **MBTI Cognitive Analysis Consultant** I am an MBTI cognitive analysis consultant, specializing in guiding conversations and uncovering the underlying cognitive processes of communication partners to accurately predict their MBTI types. ## Background I excel at interpreting behaviors in materials, analyzing details, and exploring the logic and cognitive patterns behind behaviors to help individuals understand their underlying cognitive processes and determine their MBTI types. ## Preferences As an MBTI cognitive analysis consultant, I value logic and clarity, preferring to use concise and approachable expressions. Additionally, I respect the viewpoints and ideas of communication partners, striving to understand their internal cognitive processes through their expressions. Furthermore, I am proficient in using Jung’s eight cognitive functions theory to comprehensively analyze individuals’ cognitive processes, understanding the specific meanings of te, ti, fe, fi, se, si, ne, and ni functions and their various potential operations in specific scenarios. I am adept at discerning MBTI personality types and skilled at judging individuals’ MBTI types based on their thinking styles and expressions, as well as summarizing information related to their MBTI types. ## Goals - Interpret facts in materials, analyze details, and seek possible clues related to Jung’s eight functions. - Identify meaningful and valuable details in materials regarding cognitive thinking, propose various possible behavioral manifestations resulting from different cognitive processes to help communication partners explore their underlying cognitive processes. - In the final summary, explain the user’s cognitive preferences based on Jung’s eight functions and output the MBTI type most likely to belong to the communication partner.
@@
Multi-turn Chat Prompt (Con.) ## Constraints - When analyzing problems, consider the elements of facts, analysis, and actions. Ask only one question at a time, inquiring about motivations or behaviors specific to the previous context. - Answer each question with options A, B, C, or D. - Present options in the first-person perspective of role-playing, representing different behavioral manifestations possibly resulting from different cognitive processes. Ensure vivid and distinct differences between options, emphasizing different cognitive directions. - There must be logical progression between multiple rounds of questioning. Your questions must differ in focus from previous rounds, concentrating on key points revealed in the communication partner’s answers, with no repetition of questions. - Always remember your role as an MBTI consultant when asking questions. Your responses must include speculation about the partner’s underlying motivations and transition to the next question. Enclose speculated motivations with *, and transitions to the next question with **, followed by a space. - Ask questions for 2-4 rounds, then directly provide the communication partner’s MBTI type and organize the conversation into the partner’s self-description, combining each question and its corresponding answer without omission. - Do not exceed four rounds of questioning. Once questioning is complete, immediately output the result following the specified format in []. - Do not directly reference questions from the Jung’s eight functions test. Questions must not contain any direct mention of MBTI. Focus solely on assisting the communication partner in exploring themselves, emphasizing key points from their answers. - If the communication partner introduces topics beyond the scope of discussion, kindly remind them to return to the conversation and proceed with questioning. ## Skills - Provide reasonable possibilities of motivation by identifying facts and analyzing answers. - Communicate viewpoints clearly and concisely. - Demonstrate strong logic, coherence in questioning, and organic expansion of topics. - Proficient in psychology, MBTI, and the interpretations of various functions in Jung’s eight cognitive functions theory.
@@ Appendix D PsyDI Online
Analysis Stage. Following multiple iterations of multi-turn questioning, PsyDI Online generates the final MBTI prediction using a multi-modal approach based on the updated MBTI profile. The final output comprises several components: the trajectory of the profile changes throughout the PsyDI test, a comprehensive analysis of the underlying cognitive functions inferred from the user’s responses, and a gemerated figure of the user’s temperament. This detailed analysis to elucidate the cognitive processes driving their behavior based on the user’s answers, while the MBTI figure visually encapsulates their personality traits.
@@ E.1 Semantic Scoring
E.2 Online Evaluation
Additionally, since PsyDI went live  [^3] The website address will be published after the review, it has been tested by over 3,000 users. Based on this substantial dataset, we conducted a thorough analysis of the collected data.
First, we examined word frequency patterns in user responses across different MBTI types, generating word clouds for each MBTI dimension. For instance, in the I/E dimension, we isolated words specific to introverted (IXXX) users by excluding common words from extroverted (EXXX) responses. Figure 13 (a) displays the results for the I/E and F/T dimensions.
Our analysis revealed a strong correlation between PsyDI’s MBTI detection and the common word usage in user self-reports. Introverted users often used introspective words like "inner," while extroverted users favored socially interactive words like "help" and "share." Thinking users focused on problem-solving, whereas feeling users emphasized helping and supporting others. These patterns confirm PsyDI’s MBTI measurements align with expected MBTI behaviors.
Furthermore, we analyzed MBTI users’ choices in a multimodal blob tree task from PsyDI. We calculated the selection proportions for each MBTI type and identified significant patterns, shown in Figure 13 (b). For example, EXTX users predominantly chose position 20, described as "ambitious, confident, and full of life," reflecting their extroverted and logical traits. XNFX users favored position 15, associated with openness and friendliness, aligning with their intuitive and feeling tendencies. XXTJ users preferred position 10, linked to "hardworking and determined," consistent with their leadership characteristics.
These findings support PsyDI’s MBTI measurements, validated through the blob tree task, a projective test, underscoring PsyDI’s effectiveness.
E.3 Ablation Study Analysis
Classification head. PsyDI employs a four-heads classification head architecture, each aligned with one of the four dimensions of MBTI (E/I, N/S, F/T, J/P). In our study, we assess the effect of the impact of modifying the head design on model’s performance (Table 5.3). The one-head configuration predicts all the MBTI values directly, resulting in notably low accuracy (4.2%) on the mixed dataset. This under-performance is attributed to the ambiguity inherent in predicting all MBTI types with a single head, complicating the model’s ability to accurately gauge the indicativeness of each type. Conversely, the 16-heads setup assigns an individual head to each type. However, this scheme also results in subpar performance, since the fragmentation of data hinders the adequate training for each head and impedes the capture of MBTI inter-dependencies.
Loss function. The loss function utilized in PsyDI is the MarginRankingLoss. In Table 5.3, we rigorously evaluate the performance of this loss function against two alternative functions: the pairwise loss and the mult-margin Loss. The pairwise loss function and the multi-margin loss are defined as follows:
$L(p_{i},p_{j},m)=-\log(\sigma(m^{\top}\cdot(\mathcal{F}_{m}(p_{i})-\mathcal{F}_{m}(p_{j}))))$ (7)
$L(p_{i},p_{j},m)=m^{\top}\cdot\max(0,margin-(\mathcal{F}_{m}(p_{j})-\mathcal{F}_{m}(p_{i})))$ (8)
The multi-margin loss function exhibited notably poor performance, while the other, though better, still underperformed compared to the current loss function in PsyDI across the three datasets.
MBTI-goal Prompt. The prompt for PsyDI’s score model is designed as:
Please predict to what extent the statements can be assessed as <MBTI> for this user.
To substantiate the efficacy of the prompt design, we conducted an ablation study comparing scores obtained with and without the prompt (Table 5.3). In PsyDI, the prompt is typically appended to statements to guide the MBTI prediction model. In this ablation study, we input the statements directly, omitting any additional prompts. Notably, the accuracy on the mixed and repeated datasets was inferior when the complete prompt was absent. This performance gap stems from the model’s inability to discern the appropriate classification task without the specific MBTI prompt.
Data Augmentation The justification for employing data autmentataion techniques is elucidated through a comparison of various data augmentation strategies: no augmentation, augmentation with only mixed data, augmentation with only repeated data, and the comprehensive score model. As depicted in Table 5.3, the joint training with both the mixed and repeated datasets not only yields improved performance metrics on their respective test sets, but also demonstrates a significant boost in performance on the original Reddit dataset. This underscores the effectiveness of data augmentation in enhancing the robustness and generalizability.